OpenStudy (anonymous):

Solve and type in ordered pair. x+9y=6 x=9y-5 I would change the x=9y-5 into the same form as x+9y=6 so x-9y=5 right? Then what do I do?

OpenStudy (anonymous):

Can I solve and get the coordinates by elimination process?

OpenStudy (anonymous):

that's right, use elimination

OpenStudy (anonymous):

Ok thank you

OpenStudy (anonymous):

How would that work with x and 9y's being in the same spots? Which one would I eliminate?

OpenStudy (anonymous):

f(x) = −2x + 8, find x if f(x) = 22 -2x+8 = 22 -2x = 22-8 -2x = 14 x = -7

OpenStudy (anonymous):

that's for ur other question

OpenStudy (anonymous):

x+9y=6 x-9y=5 x=1?

OpenStudy (anonymous):

the equation should be: x+9y=6 x-9y = -5

OpenStudy (anonymous):

Oh ok so the answer would be x=1? right?

OpenStudy (anonymous):

hmm, how did u get it? show me

OpenStudy (anonymous):

I did x+9y=6 and then added x-9y=-5 and canceled out the 9y-9y, and then 6-5 is one leaving me with x=1.

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

ok, bad at math, we can do another way, which is subtitution. we have: x+9y=6 x=9y-5 then just plug x=9y-5 into x+9y=6

OpenStudy (anonymous):

I'm sorry I just became very confused

OpenStudy (anonymous):

I did x+9y=6 and then added x-9y=-5 and canceled out the 9y-9y, and then 6-5 = 1, till this u are right, but u forget to do x+x = 2x 2x=1 x =1/2

OpenStudy (anonymous):

OOOHHHH ok

OpenStudy (anonymous):

So then to find y I replace x with 1/2? right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so I do that for both equations too right?

OpenStudy (anonymous):

no need, do it just to 1 equation

OpenStudy (anonymous):

ok what do I do with the y in the equation I do? Sorry I am stupid at this

OpenStudy (anonymous):

x = 1/2, plug into x+9y=6

OpenStudy (anonymous):

ok so that would be 1/2+9y=6

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok than would I subtract/add 9y to both sides?

OpenStudy (anonymous):

9y = 6 - 1/2, got it?

OpenStudy (anonymous):

I think, so to find y I would make the equation like this 2*1/2-9y=6 1-9y=6 -9y=5 y=-.6?

OpenStudy (anonymous):

is that even close to right?

OpenStudy (anonymous):

2*1/2-9y=6 ---> it's mean 2x-9y = 6 --->but our equation is x+9y=6 my answer : 1/2 +9y=6 9y = 11/2 y = 11/18 (x,y) = (1/2 , 11/18) to check ur question is correct or no, plug the order pair into initial equation: x+9y=6 1/2 + 9(11/18) = 1/2 + 11/2 = 12/2 = 6

OpenStudy (anonymous):

ok, thank you so much for help